# Collaborative and parametric insurance on the Ethereum blockchain

### v2

@@ 5 Interaction with the smart contract
To test the smart contract, three Ethereum accounts were created to represent the roles of the smart contract owner, surplus provider, and policyholder. The addresses of these accounts (plus that of the smart contract) are provided in Table 3.
@@
Our smart contract allows for the underwriting of parametric insurance contracts to compensate policyholders if the height of precipitation exceeds a threshold $\bar{Q}=5$ on a particular day in the year 2025 in either Marseille or Strasbourg. The Cornish-Fisher approximation is employed as soon as we reach 5 model points. The day of the year is represented as an integer between 1 and 365. We designed the following scenario to illustrate the functions available in the smart contract described in Section 4.
@@
The final event triggers the smart contract to file for bankruptcy, as the surplus $X$ falls below the Minimum Capital Requirement (MCR) threshold. Consequently, the premiums for policies #$\#6$, #$\#7$, #$\#8$, #$\#9$, and #$\#10$ are refunded to the policyholders. The remaining surplus is then returned to the surplus provider.
@@
Deploying the contract results in a “contract creation” transaction on the blockchain. Each subsequent function call also generates a transaction, which is recorded on the blockchain. These transactions can be retrieved and organized into a data frame, the structure of which is detailed in Table 4.
@@
hash object 0xe4$\ldots$ The unique hash value identifying the transaction
@@
blockHash object 0x0248$\ldots$ The unique hash value identifying the block that contains the transaction
@@
from object 0xe8e7$\ldots$ The address of the sender who initiated the transaction
to object 0xbe03$\ldots$ The address of the recipient of the transaction
@@
input object 0x6080$\ldots$ The data field containing information necessary for the transaction
contractAddress object 0xbe03$\ldots$ The address of the contract created as a result of the transaction
@@
methodId object 0x6080$\ldots$ The identifier for the method being called in a smart contract
@@
[…]
@@
In addition to the list of all transactions, we can retrieve the log of events emitted by the smart contract. A common practice is to associate a specific event with each function. Function calls impact the state of the system, and events provide a mechanism to monitor state transitions effectively. However, care must be taken when defining events, as events that contain a large amount of information will require more data storage and, consequently, increase transaction costs. The definition of the events are provided in Table 5.
@@
customer__\__address hex $0x456\ldots$ Ethereum address
@@
customer__\__address hex $0x456\ldots$ Ethereum address
@@
The events used in this study are defined in the Solidity code, which is included in the online supplementary materials accompanying this article. Using Python, we retrieve the event logs to track participant balances over time. Figure 9 illustrates the balances, in ETH and before transaction fees, of the Ethereum addresses associated with the Owner, the Surplus Provider, the Policyholder, and the Smart Contract after each function call.
@@
We observe that the scenario was particularly beneficial to the Policyholder, who received three compensation payments and a refund for five policies. The event logs also enable us to visualize the trajectories of the various processes defined in Section 3 and Section 4. Specifically, Figure 10 displays the trajectories of $X$ (surplus), $B$ (balance), SCR (Solvency Capital Requirement), and MCR (Minimum Capital Requirement) as they evolved throughout the scenario.
@@
Note the difference between the surplus and the balance of the smart contract: the balance of the smart contract increases at each "Underwrite" event, while the surplus remains constant. Until contract #$\#5$ is underwritten, the Solvency Capital Requirement (SCR) and Minimum Capital Requirement (MCR) are exactly equal to the sum of all potential compensations to be paid. Once contract #$\#5$ is underwritten, the number of model points reaches $5$, and the solvency capitals are calculated via the Cornish-Fisher approximation, causing a drop in the value of the solvency capitals. After the settlement of contract #$\#3$, the exchange rate of protocol tokens against ETH appreciates, which is why the surplus provider decides to cash out part of her investment. The compensations paid after the settlements of contracts #$\#1$, #$\#2$, and #$\#4$, combined with the fact that the settlement of contract #$\#5$ reduces the number of model points below $5$, cause the smart contract to reset as the surplus becomes smaller than the MCR. The status of all active policies changes to "cancelled." Premiums are refunded to the policyholders, and investors receive any remaining balance from the smart contract.
@@
The drop in value of both the Minimum Capital Requirement (MCR) and the Solvency Capital Requirement (SCR) after including contract ##\##5 might seem surprising at first glance. Under Solvency II regulations, there is an absolute floor on the MCR value, set at least to €2.7 million euros for non-life insurance companies. We have proposed an alternative method for defining this floor, based on the sum of exposures when the number of independent risks is too small (fewer than 5 in Figure 10 for illustration purposes, but typically fewer than hundreds or thousands in real-life scenarios). In practice, a decrease in solvency capital can occur during the transition from the standard formula to internal models, which somewhat corresponds to adopting normal approximations in our example.
@@ 6 Limitations and perspectives
Our smart contract is exposed to the risk of a "mass investor lapse," where investors may collectively withdraw their tokens. Stronger incentive mechanisms could be implemented to encourage investors to retain their tokens, particularly when the contract’s balance approaches the risk capital threshold. One potential approach is to introduce a coin age principle, where tokens held for longer periods gain additional value. This concept has already been explored in the implementation of the cryptocurrency PeerCoin [^9] See https://www.peercoin.net/. A traditional approach to managing the surplus of an insurance company is to implement a dividend policy that rewards investors or, more generally, token holders. This can be achieved using classical optimal dividend strategies. Barrier or band strategies, for instance, have been shown to be optimal across various modeling frameworks (see the survey by Avanzi [3]). In our framework, we operate within a discrete-time risk model with capital injections and withdrawals. This approach is consistent with the structure of our system, where wealth is only inspected at discrete points—namely, during settlements or when an investor funds or withdraws from the contract.
@@
In our approach, surplus management is handled through proxies referred to as the Minimum Capital Requirement (MCR) and Solvency Capital Requirement (SCR). If such a smart contract were to be classified as an insurance product, it would need to comply with the absolute minimum MCR threshold prescribed by Solvency II in Europe. This requirement could lead to a chicken-and-egg problem, as achieving the necessary critical mass of funds would be essential to meet the MCR. Additionally, calculating the SCR and MCR would typically involve integrating multiple risk modules, adding further complexity and thus incur additional implementation costs.
@@
If the contract is not classified as an insurance product, it could lead to tax and regulatory implications. To ensure recognition as insurance, it is crucial to keep basis risk sufficiently low. Effective basis risk management not only protects policyholders but also mitigates the risk of the contract being reclassified as gambling. In some countries, policyholders may be required to provide brief proof of loss to validate claims. While we assume that the occurrence of a compensation-triggering event inherently results in a loss for the policyholder, real-world scenarios might necessitate additional efforts from either the policyholder or the smart contract’s governance system to address such requirements. Basis risk management is a central problem when dealing with parametric insurance solution and as such under investigation in many recent research work like that of Niakh et al. [26].
@@
In conclusion, we have developed a modest beta version of a smart contract for parametric insurance on the Ethereum blockchain. The Solidity code for the smart contract and the Python scripts for analyzing blockchain data are freely available in the smart__\__parametric__\__insurance GitHub repository. Beyond serving as a proof of concept, this work identifies several research questions that need to be addressed to make the product fully operational. We hope that our open-source beta version will provide a foundation for other researchers to build upon, facilitating the development of enhanced features and overcoming the limitations we have outlined.

### v3

@@ 5 Interaction with the smart contract
To test the smart contract, three Ethereum accounts were created to represent the roles of the smart contract owner, surplus provider, and policyholder. The addresses of these accounts (plus that of the smart contract) are provided in Table˜3.
@@
Our smart contract allows for the underwriting of parametric insurance contracts to compensate policyholders if the height of precipitation exceeds a threshold $\bar{Q}=5$ on a particular day in the year 2025 in either Marseille or Strasbourg. The Cornish-Fisher approximation is employed as soon as we reach 5 model points. The day of the year is represented as an integer between 1 and 365. We designed the following scenario to illustrate the functions available in the smart contract described in Section˜4.
@@
The final event triggers the smart contract to file for bankruptcy, as the surplus $X$ falls below the Minimum Capital Requirement (MCR) threshold. Consequently, the premiums for policies #6\#6, #7\#7, #8\#8, #9\#9, and #10\#10 are refunded to the policyholders. The remaining surplus is then returned to the surplus provider.
@@
Deploying the contract results in a “contract creation” transaction on the blockchain. Each subsequent function call also generates a transaction, which is recorded on the blockchain. These transactions can be retrieved and organized into a data frame, the structure of which is detailed in Table˜4.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\texttt{gasPrice}\times\texttt{gasUsed}.$
The value of a unit of gas (gasPrice) fluctuates over time, as it depends on the level of activity on the Ethereum network. High network congestion leads to higher gas prices. Another factor influencing transaction costs is the amount of data being recorded, which is reflected in the input field of the transaction details. The "contract creation" transaction is by far the most expensive, with a cost of approximately $\text{ETH}0.035$, due to the significant amount of data required to deploy the smart contract. In contrast, subsequent transactions typically cost between $\text{ETH}0.002$ and $\text{ETH}0.008$, as illustrated in Figure˜8.
@@
In addition to the list of all transactions, we can retrieve the log of events emitted by the smart contract. A common practice is to associate a specific event with each function. Function calls impact the state of the system, and events provide a mechanism to monitor state transitions effectively. However, care must be taken when defining events, as events that contain a large amount of information will require more data storage and, consequently, increase transaction costs. The definition of the events are provided in Table˜5.
@@
customer_\_address hex $0x456\ldots$ Ethereum address
@@
customer_\_address hex $0x456\ldots$ Ethereum address
@@
The events used in this study are defined in the Solidity code, which is included in the online supplementary materials accompanying this article. Using Python, we retrieve the event logs to track participant balances over time. Figure˜9 illustrates the balances, in ETH and before transaction fees, of the Ethereum addresses associated with the Owner, the Surplus Provider, the Policyholder, and the Smart Contract after each function call.
@@
We observe that the scenario was particularly beneficial to the Policyholder, who received three compensation payments and a refund for five policies. The event logs also enable us to visualize the trajectories of the various processes defined in Section˜3 and Section˜4. Specifically, Figure˜10 displays the trajectories of $X$ (surplus), $B$ (balance), SCR (Solvency Capital Requirement), and MCR (Minimum Capital Requirement) as they evolved throughout the scenario.
@@
Note the difference between the surplus and the balance of the smart contract: the balance of the smart contract increases at each "Underwrite" event, while the surplus remains constant. Until contract #5\#5 is underwritten, the SCR and MCR are exactly equal to the sum of all potential compensations to be paid. Once contract #5\#5 is underwritten, the number of model points reaches $5$, and the solvency capitals are calculated via the Cornish-Fisher approximation, causing a drop in the value of the solvency capitals. After the settlement of contract #3\#3, the exchange rate of protocol tokens against ETH appreciates, which is why the surplus provider decides to cash out part of her investment. The compensations paid after the settlements of contracts #1\#1, #2\#2, and #4\#4, combined with the fact that the settlement of contract #5\#5 reduces the number of model points below $5$, cause the smart contract to reset as the surplus becomes smaller than the MCR. The status of all active policies changes to "cancelled." Premiums are refunded to the policyholders, and investors receive any remaining balance from the smart contract.
@@
The drop in value of both the MCR and the SCR after including contract #\#5 might seem surprising at first glance. Under Solvency II regulations, there is an absolute floor on the MCR value, set at least to €2.7 million euros for non-life insurance companies. We have proposed an alternative method for defining this floor, based on the sum of exposures when the number of independent risks is too small (fewer than 5 in Figure˜10 for illustration purposes, but typically fewer than hundreds or thousands in real-life scenarios). In practice, a decrease in solvency capital can occur during the transition from the standard formula to internal models, which somewhat corresponds to adopting normal approximations in our example.
The Solidity code for the smart contract and the Python scripts for analyzing blockchain data are freely available in the smart_\_parametric_\_insurance GitHub repository.
@@ 6 Limitations and perspectives
Our smart contract is exposed to the risk of a "mass investor lapse," where investors may collectively withdraw their tokens. Stronger incentive mechanisms could be implemented to encourage investors to retain their tokens, particularly when the contract’s balance approaches the risk capital threshold. One potential approach is to introduce a coin age principle, where tokens held for longer periods gain additional value. This concept has already been explored in the implementation of the cryptocurrency PeerCoin [^9] See https://www.peercoin.net/. A traditional approach to managing the surplus of an insurance company is to implement a dividend policy that rewards investors. This can be achieved using classical optimal dividend strategies. Barrier or band strategies, for instance, have been shown to be optimal across various modeling frameworks (see the survey by Avanzi [3]). In our framework, we operate within a discrete-time risk model with capital injections and withdrawals. This approach is consistent with the structure of our system, where wealth is only inspected at discrete points—namely, during settlements or when an investor funds or withdraws from the contract.
@@
In our approach, surplus management is handled through proxies referred to as the Minimum Capital Requirement (MCR) and Solvency Capital Requirement (SCR). If such a smart contract were to be classified as an insurance product, it would need to comply with the absolute minimum MCR threshold prescribed by Solvency II in Europe.
@@
If the contract is not classified as an insurance product, it could lead to tax and regulatory implications. To ensure recognition as insurance, it is crucial to keep basis risk sufficiently low. Effective basis risk management not only protects policyholders but also mitigates the risk of the contract being reclassified as gambling. In some countries, policyholders may be required to provide a brief proof of loss to validate claims. While we assume that the occurrence of a compensation-triggering event inherently results in a loss for the policyholder, real-world scenarios might necessitate additional efforts from either the policyholder or the smart contract’s governance system to address such requirements. Basis risk management is a central problem when dealing with parametric insurance solution and as such under investigation in many recent research work like that of Niakh et al. [26].
@@
In conclusion, we have developed a modest beta version of a smart contract for parametric insurance on the Ethereum blockchain. The Solidity code for the smart contract and the Python scripts for analyzing blockchain data are freely available in the smart_\_parametric_\_insurance GitHub repository. Beyond serving as a proof of concept, this work identifies several research questions that need to be addressed to make the product fully operational. We hope that our open-source beta version will provide a foundation for other researchers to build upon, facilitating the development of enhanced features and overcoming the limitations we have outlined.